# Causally-Guided Pairwise Transformer - Towards Foundational Digital Twins in Process Industry

### v1

@@ 1.3 Limitations of Existing Time-Series modelling Paradigms
Figure 1: (a) Channel-Independent (CI): A shared function $f$ processes each variable independently, learning universal temporal patterns but ignoring cross-channel dependencies. (b) Channel-Dependent (CD): A single function $f$ processes all variables $(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)})$ jointly, i.e. mixed, to directly model their interdependencies. (c) CGPT (CI+CD Pairwise): A single function $f$ processes pairs of targetcontext variables ($(\mathbf{X}^{(i)}$$\mathbf{C}^{(j)})$) jointly (CD). All pairs w.r.t a specific target are processed independently (CI). The outputs from these pairwise computations are aggregated in the latent space (see Sec. 2). This captures specific dependencies between the target and one context variable (CD) while generalizing across pairs (CI). (Figure adapted from [6, 10]).
@@ 1.4 Contributions
Figure 2: An overview of the proposed model architecture, which operates in three stages. (1) A CI module first extracts a latent vector $z_{i}$ for each input time series $C_{i}(t)$ using a shared temporal transformer (CI). (2) Guided by a causal graph, pairwise interaction modules then compute influence vectors, such as $I(i\leftarrow j)$, to mix a cause node $C_{j}$ on a target child node $C_{i}$. (3) Finally, the influence vectors from all causal parents are aggregated and added to the target’s initial latent vector to form a final representation $z^{\prime}_{i}$. This representation is then passed through a prediction head to generate the forecast $\hat{y}_{i}$.
@@ 2 CGPT: Causally-Guided Pairwise Transformer - Towards Foundational Digital Twins
Our objective is to develop a function $f_{\theta}:\mathcal{I}\to\mathcal{O}$ that maps historical system states ($\mathcal{I}$) to future values of specific target variables ($\mathcal{O}$), effectively capturing intricate temporal dependencies within industrial processes. Let $\mathbf{S}_{t}\in\mathbb{R}^{N_{vars}}$ represent the multivariate state of an industrial process at discrete time step $t$. This state comprises two main categories of variables:
  • Internal Process State Measurements ($\mathbf{X}_{t}\in\mathbb{R}^{D_{X}}$): Continuously sensed data that directly describe the current physical or chemical state of the ongoing industrial process itself (e.g., temperatures inside a reactor, pressures in a pipeline, motor current, air flow). These are typically reactive measurements reflecting the process’s internal dynamics.
  • Operational and Product-Specific Variables ($\mathbf{U}_{t}\in\mathbb{R}^{D_{U}}$): Variables that characterize the operational configuration or external context of the process. This includes actively controlled parameters (e.g., machine setpoints, valve positions, energy input) as well as variables defining the specific product being manufactured or environmental conditions (e.g., product type, batch number, raw material properties, ambient humidity). These variables often dictate the desired process behavior or represent external influences.
Thus, the total number of observed variables is $N_{vars}=D_{X}+D_{U}$. The model input is a sequence of past observations over a context length $L_{ctx}$: $\mathcal{S}_{hist}=\left((\mathbf{X}_{t-L_{ctx}+1},\mathbf{U}_{t-L_{ctx}+1}),\dots,(\mathbf{X}_{t},\mathbf{U}_{t})\right)\in\mathbb{R}^{L_{ctx}\times N_{vars}}$. The model $f_{\theta}$ predicts a sequence of future target values, $\hat{\mathcal{Y}}_{future}$, over a prediction horizon $H_{pred}$: $\hat{\mathcal{Y}}_{future}=f_{\theta}(\mathcal{S}_{hist}),\quad\text{where }\mathcal{Y}_{future}=(\mathbf{Y}_{t+1},\dots,\mathbf{Y}_{t+H_{pred}})$. Here, $\mathbf{Y}_{t+h}\in\mathbb{R}^{N_{target}}$ represents the target variables (Key Performance Indicators or KPIs) at future step $h\in\{1,\dots,H_{pred}\}$, which are typically a subset of the input variates forming $\mathcal{S}_{hist}$. These KPIs can be process-related (e.g., forecasting internal temperature of a kiln) or more product-related (e.g., forecasting downstream quality measurements). The remaining variables in $\mathcal{S}_{hist}$ then serve as covariates for the prediction task.
@@ 2.1 Architectural Components
The CGPT model, denoted as $f_{\theta}$, is a deep learning architecture that operationalizes the principle of forecast decomposition. For a given pair of a target time series $\mathbf{x}^{(i)}$ and a context time series $\mathbf{x}^{(j)}$, it predicts the future of the target by explicitly modelling each context’s influence.
@@
1. Independent Temporal Feature Extraction (CI): Each time-series instance within the pair is first normalized individually using Reversible Instance Normalization (RevIN) [17] within a context window, this dampens the impact of distributional shifts. The normalized series are then segmented into smaller non-overlapping patches, a technique inspired by PatchTST [8] and commonly found in the image processing domains. These patches serve as fundamental units for subsequent processing by the deep Transformer architecture, reducing computational complexity and increasing the local receptive field. Specifically, an input time series $\mathbf{x}\in\mathbb{R}^{L_{ctx}}$ of length $L_{ctx}$ is permuted and unfolded into patches $\mathbf{P}\in\mathbb{R}^{N_{p}\times P}$, where $P$ is the patch size, and $N_{p}=\lfloor(L_{ctx}-P)/S\rfloor+1$ is the number of patches given stride $S$. This operation is handled by a dedicated patching module.
The initial token sequences for the target variable, $\mathbf{X}_{0}^{(i)}$, and the context variable, $\mathbf{C}_{0}^{(j)}$, are processed through a shared-weight Transformer encoder stack of $E_{\text{layers}}$ layers. Each variable’s sequence is processed independently, meaning self-attention is computed only within each sequence, not between them at this stage. This design choice aligns with the core principle of Channel-Independent (CI) models [8], promoting the learning of generalizable, variable-agnostic temporal feature extractors. The encoder transforms the input tokens into context-aware latent representations, denoted as $\mathbf{Z}_{E}^{(i)}$ and $\mathbf{Z}_{E}^{(j)}\in\mathbb{R}^{N_{p}\times D_{\text{model}}}$. These representations encapsulate the temporal dynamics of each series, providing the foundational features for the subsequent forecasting and influence modelling heads.
$\mathbf{Z}_{E}^{(v)}=\text{Encoder}(\mathbf{Z}_{0}^{(v)})\quad\text{for }v\in\{i,j\}$ (1)
2. Additive Pairwise Interaction modelling (CD): The latent vectors of target $i$ and context $j$ are concatenated and passed through an MLP to the model dimension. We use a simple summing operator to aggregate the latent pairwise vectors with the latent target vectors. Other, more advanced concepts may be explored, like learnable channel reconciliation strategies and more advanced channel mixing layers like pairwise cross-attention (see Sec.4).
$\hat{\mathbf{y}}^{(i)}_{\text{final}}={\mathbf{z}}^{(i)}+\sum_{k=1}^{K}{\mathbf{z}}^{(i,j_{k})}_{\text{influence}}$ (2)
Crucially, the entire model is trained end-to-end with a single loss function (e.g., MSE) on the final prediction $\hat{\mathbf{y}}^{(i)}_{\text{final}}$. The decomposition is therefore learned implicitly as a result of the architectural separation of the information flows, rather than through explicit loss terms on the individual base and influence components.
@@ 3.1 Experimental Setup
Multi-Stage Fact.2 Public 56 14088 S1…0.U.Err.
Amino Emissions3  [19] Public 67 5409 2-Amino…C4H11NO
@@
We evaluate against two strong baselines (CI and CD) and three variants of CGPT. The baselines are DLinear, a univariate model (CI) that only uses the target’s own history, and MLP-Baseline, a generic multivariate model (CD) that flattens all channel histories and processes them with a deep MLP. Our primary model, LeakyPairwise CGPT (Ours), combines CI and CD approaches by computing pairwise influences using the target’s own encoded history. We also test two ablations: StrictPairwise CGPT, where the target’s history in the pairwise interaction module is replaced with a generic, learnable embedding for the influence calculation, and PureInfluence CGPT, which predicts using only the sum of influences, making the model completely blind to the target’s history in the final step.
@@ 3.4 Experiment 2: One-Step Forecasting with CGPT vs CD and CI Baselines
For the one-step-ahead forecasting task (context=96, horizon=1), the results in Tab. 3 show a clear preference for simplicity. The DLinear model, which excels at capturing autoregressive patterns, decisively outperforms the more complex CD models. Among the remaining models, the LeakyPairwise and StrictPairwise variants also show a clear advantage over the standard MLP, performing better on most datasets.

### v2

@@ 1.3 Limitations of Existing Time-Series modelling Paradigms
Figure 1: (a) Channel-Independent (CI): A shared function $f$ processes each variable independently, learning universal temporal patterns but ignoring cross-channel dependencies. (b) Channel-Dependent (CD): A single function $f$ processes all variables $(\mathbf{X}^{(1)},\dots,\mathbf{X}^{(N)})$ jointly, i.e. mixed, to directly model their interdependencies. (c) CGPT (CI+CD Pairwise): A single function $f$ processes pairs of targetcontext variables ($(\mathbf{X}^{(i)}$$\mathbf{C}^{(j)})$) jointly (CD). All pairs w.r.t a specific target are processed independently (CI). The outputs from these pairwise computations are aggregated in the latent space (see Sec. 2). This captures specific dependencies between the target and one context variable (CD) while generalizing across pairs (CI). (Figure adapted from [6]).
@@ 2 CGPT: Causally-Guided Pairwise Transformer - Towards Foundational Digital Twins
Our objective is to develop a function $f_{\theta}$ that maps the historical state of a multivariate industrial process to the future value of a specific target variable, effectively capturing intricate, causally-informed temporal dependencies. Let $\mathbf{C}_{t}=(C_{1}(t),\dots,C_{N_{vars}}(t))\in\mathbb{R}^{N_{vars}}$ represent the multivariate state of the process at discrete time step $t$. The individual channels $C_{n}(t)$ in this state vector can be broadly categorized:
  • Internal Process State Measurements: Continuously sensed data that directly describe the physical or chemical state of the process (e.g., temperatures inside a reactor, pressures in a pipeline, motor current, air flow). These are typically reactive measurements reflecting the process’s internal dynamics.
  • Operational and Product-Specific Variables: Variables characterizing the operational configuration or external context. This includes actively controlled parameters (e.g., machine setpoints, valve positions) and variables defining the product or environment (e.g., product type, raw material properties, ambient humidity). These channels often act as causal drivers for the process behavior.
Figure 2: An overview of the proposed model architecture, which operates in three stages. (1) A CI module first extracts a latent vector $z_{i}$ for each input time series $C_{i}(t)$ using a shared temporal transformer. (2) Guided by a causal graph, pairwise mixing modules compute a vector $I(i,j)$ by mixing the latent representations of a cause node ($C_{j}$) and the target node ($C_{i}$). (3) Finally, these mixed vectors from all causal parents are aggregated into $z_{\text{int}}$ and added to the target’s initial latent vector to form a final representation $z^{\prime}_{i}$. This representation is then passed through a prediction head to generate the forecast $\hat{y}_{i}$.
The model’s task is to predict the future of a chosen target channel, $C_{i}$, based on its own history and the history of a pre-defined set of causal context channels, $\{C_{j}\}_{j\in\mathcal{J}}$. The input to the model $f_{\theta}$ is a sequence of past observations for these relevant channels over a context length $L_{ctx}$. The model then outputs a forecast for the target channel, $\hat{y}_{i}$, over a given prediction horizon. The target variable $C_{i}$ is typically a Key Performance Indicator (KPI), such as a critical process temperature or a downstream quality measurement. The context channels in $\mathcal{J}$ serve as covariates for the prediction task, providing the necessary information to model the system’s dynamics accurately.
@@ 2.1 Architectural Components
The CGPT model, denoted as $f_{\theta}$, operationalizes this predictive task in three distinct stages, as illustrated in Fig. 2.
@@
1. Independent Temporal Feature Extraction (CI): Each input time series channel is first normalized individually using Reversible Instance Normalization (RevIN) [17] within a context window; this dampens the impact of distributional shifts. The normalized series are then segmented into smaller non-overlapping patches, a technique inspired by PatchTST [8] and commonly found in image processing domains. These patches serve as fundamental units for subsequent processing by the deep Transformer architecture. Specifically, an input time series $C_{n}\in\mathbb{R}^{L_{ctx}}$ is unfolded into a sequence of patches $\mathbf{C}_{0}^{(n)}\in\mathbb{R}^{N_{p}\times P}$, where $P$ is the patch size and $N_{p}$ is the number of patches. The initial token sequences for each channel are processed through a shared-weight Transformer encoder stack. Each channel’s sequence is processed independently, meaning self-attention is computed only within each sequence. This design choice aligns with the core principle of Channel-Independent (CI) models [8], promoting the learning of generalizable temporal features. The encoder transforms the input tokens of each channel $C_{n}$ into a latent vector, denoted as $z_{n}$. These vectors encapsulate the temporal dynamics of each series, providing the basis for the subsequent interaction modeling.
$\mathbf{z}_{n}=\text{Encoder}(\mathbf{C}_{0}^{(n)})\quad\text{for }n\in\{1,\dots,N_{vars}\}$ (1)
2. Pairwise Interaction Modeling (CD): This stage models how information from a set of causal parents, indexed by $\mathcal{J}$, is mixed with that of the target variable $C_{i}$. The core CD mechanism is a parameter-shared pairwise CD mixing module. It computes an interaction vector for each pair, $\mathbf{I}(i,j)$, by applying a feed-forward network (FFN) to the concatenation (denoted as $\cdot$$\cdot$]) of the target and parent latent vectors. 3. Additive Aggregation and Prediction: The individual interaction vectors, $\mathbf{I}(i,j)$, from all causal parents are aggregated via summation, denoted as $\mathbf{z}_{\text{int}}$ in Fig. 2. This total interaction component is then additively combined with the target’s original latent vector $\mathbf{z}_{i}$ to form the final representation, $\mathbf{z}^{\prime}_{i}$. The entire process is captured in a single end-to-end model:
$\hat{y}_{i}=\text{PredictionHead}\left(\mathbf{z}_{i}+\sum_{j\in\mathcal{J}}\mathbf{I}(i,j)\right)$ (2)
We use a simple summing operator to aggregate these individual mixed vectors. While this approach is straightforward and effective, more advanced concepts may be explored, such as learnable channel reconciliation strategies or more advanced mixing layers like pairwise cross-attention (see Sec. 4). Crucially, the entire model is trained end-to-end using a single loss function (e.g., MSE) on the final forecast $\hat{y}_{i}$. The decomposition is therefore learned implicitly through the architectural separation of information flows.
@@ 3.1 Experimental Setup
Kaggle Multi-Stage Fact.2 Public 56 14088 S1…0.U.Err.
Carbon Amino Emissions3  [19] Public 67 5409 2-Amino…C4H11NO
@@
We evaluate against two strong baselines (CI and CD) and three variants of CGPT. The baselines are DLinear, a univariate model (CI) that only uses the target’s own history, and MLP-Baseline, a generic multivariate model (CD) that flattens all channel histories and processes them with a deep MLP. Our primary model, LeakyPairwise CGPT (Ours), combines CI and CD approaches by computing pairwise interactions using the target’s own encoded history. We also test two ablations: StrictPairwise CGPT, where the target’s history in the pairwise interaction module is replaced with a generic, learnable embedding for the interaction calculation, and PureInfluence CGPT, which predicts using only the sum of the resulting interaction vectors, making the model completely blind to the target’s history in the final step.
@@ 3.4 Experiment 2: One-Step Forecasting with CGPT vs CD and CI Baselines
For the one-step-ahead forecasting task (context=96, horizon=1), the results in Tab. 3 show a clear preference for simplicity. The DLinear model, which excels at capturing autoregressive patterns, decisively outperforms the more complex CD models on most datasets. Among the remaining models, the LeakyPairwise and StrictPairwise variants also show a clear advantage over the standard MLP, performing better on most datasets.